# Remote State Estimation over Unreliable Channels with Unreliable Feedback: Fundamental Limits

### v1

@@ V Conclusions
$\qquad\qquad+\sum_{t^{\prime}=1}^{t}\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\boldsymbol{\sigma}_{0:t^{\prime}-1}=0,\sigma_{t^{\prime}}=1\Big{)}=1$ (44)
for any $t\in\mathbb{N}_{[0,N]}$. Applying the law of total expectation for the terms $\operatorname{\mathsf{E}}[\alpha_{k}\delta_{k}]$ and $\operatorname{\mathsf{E}}[\tilde{e}_{k}^{T}\tilde{e}_{k}]$ in $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ on a partition provided by the identity (Proof:) for $t=k-1$, repeating this procedure for $k\in\mathbb{N}_{[1,N]}$, and using the definition of $\Psi^{k+1,N}_{\epsilon}(\tilde{e}_{k+1})$ along with the relation in (Proof:), we obtain (LABEL:eq:lemma:Omega). This completes the proof.
In light of Lemma 1 and from the definition of the loss function $\Phi(\epsilon,\delta)$, we can write
$\Phi(\epsilon,\delta)$ $=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\Big{[}\alpha_{k}\delta_{k}+\hat{e}_{k}^{T}\hat{e}_{k}\Big{]}$
$=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\Big{[}\alpha_{k}\delta_{k}+\tilde{e}_{k}^{T}\tilde{e}_{k}+\operatorname{tr}Q_{k}\Big{]}$
where $\delta$ is chosen based on $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, and in the second equality we used the tower property of conditional expectations. Define the loss function $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ as
$\Psi^{N}_{\epsilon}(\tilde{e}_{0}):=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\Big{[}\alpha_{k}\delta_{k}+\tilde{e}_{k}^{T}\tilde{e}_{k}\Big{]}.$
Following the fact that $\operatorname{tr}Q_{k}$ is independent of $\epsilon$ and $\delta$, to optimize $\Phi(\epsilon,\delta)$ over $\epsilon\in\mathcal{E}$, it suffices to optimize over $\epsilon\in\mathcal{E}$. In addition, observe that from the law of total probability, we have
$\qquad\qquad+\sum_{t^{\prime}=1}^{t}\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\boldsymbol{\sigma}_{0:t^{\prime}-1}=0,\sigma_{t^{\prime}}=1\Big{)}=1$ (45)
for any $t\in\mathbb{N}_{[0,N]}$. Applying the law of total expectation for the terms $\operatorname{\mathsf{E}}[\alpha_{k}\delta_{k}]$ and $\operatorname{\mathsf{E}}[\tilde{e}_{k}^{T}\tilde{e}_{k}]$ in $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ on a partition provided by the identity (Proof:) for $t=k-1$, repeating this procedure for $k\in\mathbb{N}_{[1,N]}$, and using the definition of $\Psi^{k,N}_{\epsilon}(\tilde{e}_{k})$, we can obtain (8). ∎
@@
Suppose that $(\epsilon^{o},\delta^{o})$ is a coding policy profile that belongs to the set of globally optimal solutions, which obviously cannot be empty. We prove global optimality of the policy profile $(\epsilon^{\star},\delta^{\star})$ in the claim by showing that $\Phi(\epsilon^{\star},\delta^{\star})$ cannot be greater than $\Phi^{o}:=\Phi(\epsilon^{o},\delta^{o})$. Without loss of generality, we assume that $m_{0}=0$. Similar arguments can be made for $m_{0}\neq 0$ following a coordinate transformation. The proof consists of three steps.
@@
Step 1. We say a scheduling policy is innovation-based if, at each time $k$, it depends on $\boldsymbol{\nu}_{0:k}$ instead of $\boldsymbol{y}_{0:k}$ and $\boldsymbol{z}^{b}_{0:k}$, i.e., if it can be expressed in the form $\operatorname{\mathsf{p}}(\delta_{k}|\boldsymbol{\nu}_{0:k},\boldsymbol{\delta}_{0:k-1},\boldsymbol{\theta}_{0:k-1})$. In the first step of our proof, we show that $\Phi(\epsilon^{n},\delta^{o})=\Phi(\epsilon^{o},\delta^{o})$, where $\epsilon^{n}$ is an innovation-based scheduling policy. From the definition of the innovation, (20), and (21), we can write $\boldsymbol{y}_{0:k}=\boldsymbol{\nu}_{0:k}+G_{k}\check{\boldsymbol{x}}_{0:k-1}$ and $\check{\boldsymbol{x}}_{0:k}=H_{k}\boldsymbol{\nu}_{0:k}$, where $G_{k}$ and $H_{k}$ are matrices of proper dimensions. Putting these equations together, we find $\boldsymbol{y}_{0:k}=\boldsymbol{\nu}_{0:k}+G_{k}H_{k-1}\boldsymbol{\nu}_{0:k-1}$. Therefore, $\operatorname{\mathsf{p}}_{\epsilon^{o}}(\delta_{k}|\boldsymbol{y}_{0:k},\boldsymbol{z}^{b}_{0:k},\boldsymbol{\delta}_{0:k-1},\boldsymbol{\theta}_{0:k-1})$ can equivalently be written as $\operatorname{\mathsf{p}}_{\epsilon^{n}}(\delta_{k}|\boldsymbol{\nu}_{0:k},\boldsymbol{z}^{b}_{0:k},\boldsymbol{\delta}_{0:k-1},\boldsymbol{\theta}_{0:k-1})$. This establishes that $\Phi(\epsilon^{n},\delta^{o})=\Phi(\epsilon^{o},\delta^{o})$. As our subsequent analysis is valid for any values of $\boldsymbol{z}^{b}_{0:k}$ and $\boldsymbol{\theta}_{0:k-1}$, for brevity, hereafter we omit the dependency of $\epsilon^{n}$ on these variables.
@@
Step 2. Let $\mathcal{B}(r)$ be a ball of radius $r$ centered at the origin and of proper dimension. Define $\varpi_{k}:=T_{k}\boldsymbol{\nu}_{0:k}\in\mathcal{N}$ for a specific matrix $T_{k}$, and $\sigma_{k}:=\delta_{k}\gamma_{k}$, $l_{\epsilon^{s}}(\varpi_{k}):=(\lambda+\lambda^{c}\operatorname{\mathsf{p}}_{\epsilon^{s}}(\delta_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0))\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$, and $l_{\epsilon^{n}}(\varpi_{k}):=(\lambda+\lambda^{c}\operatorname{\mathsf{p}}_{\epsilon^{n}}(\delta_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0))\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$. In the second step of our proof, we show that $\Phi(\epsilon^{s},\delta^{o})\leq\Phi(\epsilon^{n},\delta^{o})$, where $\epsilon^{s}$ is a special form of $\epsilon^{n}$ that is symmetric with respect to $\boldsymbol{\nu}_{0:k}$ at each time $k$ and such that the following conditions are satisfied:
[…]
@@
for all $r\geq 0$ with $l_{\epsilon^{s}}(\varpi_{k})$ as a radially symmetric function of $\varpi_{k}$. We know that $\sigma_{k}=0$ either when $\delta_{k}=1\;\wedge\;\gamma_{k}=0$ or when $\delta_{k}=0$ regardless of $\gamma_{k}$. Accordingly, we can write
@@
$\ =\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}=1\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0)$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}\gamma_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0,\delta_{k}=1\Big{)}$
$\qquad+\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big{)}$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}\gamma_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0,\delta_{k}=0\Big{)}$
$\ =\lambda\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}=1\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big{)}$
$\qquad+\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big{)}$
$\ =\lambda+\lambda^{c}\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big{)}$
where in the first equality we used the law of total probability and in the second equality the fact that $\gamma_{k}$ is independent of $\varpi_{k}$, $\boldsymbol{\sigma}_{0:k-1}$, and $\delta_{k}$. Define $l^{\prime}_{\epsilon^{s}}(\varpi_{k})=\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0)\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$ and $l^{\prime}_{\epsilon^{s}}(\varpi_{k})=\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0)\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$. Therefore, the above condition can be written in the following equivalent form:
[…]
for all $r\geq 0$ with $l^{\prime}_{\epsilon^{s}}(\varpi_{k})$ as a radially symmetric function of $\varpi_{k}$. By Lemma 8, to prove that $\Phi(\epsilon^{s},\delta^{o})\leq\Phi(\epsilon^{n},\delta^{o})$, it suffices to prove that $\Psi^{M}_{\epsilon^{s}}(\tilde{e}_{0})\leq\Psi^{M}_{\epsilon^{n}}(\tilde{e}_{0})$ for any $M\in\mathbb{N}_{[0,N]}$ and for any $\tilde{e}_{0}$. Observe that $\tilde{e}_{0}=K_{0}\nu_{0}$ is the same under both $\epsilon^{n}$ and $\epsilon^{s}$, and that $\delta_{0}$ has no effects on remote state estimation when the time horizon is 00. Hence, the claim holds for the time horizon is 00. We assume that the claim also holds for all time horizons from $1$ to $M-1$, and will show that the terms in $\Psi^{M}_{\epsilon^{n}}(\tilde{e}_{0})$ are not less than those in $\Psi^{M}_{\epsilon^{s}}(\tilde{e}_{0})$.
@@
×$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
@@
×$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
$=\operatorname{\mathsf{p}}_{\epsilon^{s}}\Big{(}\sigma_{k-1}=0\big{|}\boldsymbol{\sigma}_{0:k-2}=0\Big{)}$
[…]
@@
$\operatorname{\mathsf{E}}_{\epsilon^{n}}\Big{[}$ $\delta_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
$=1-\operatorname{\mathsf{p}}_{\epsilon^{n}}(\delta_{k}=0|\boldsymbol{\sigma}_{0:k-1}=0)$
@@
$=1-\operatorname{\mathsf{p}}_{\epsilon^{s}}(\delta_{k}=0|\boldsymbol{\sigma}_{0:k-1}=0)$
$=\operatorname{\mathsf{E}}_{\epsilon^{s}}\Big{[}\delta_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}.$
@@
$\tilde{e}_{t}$ $=A_{t-1}\tilde{e}_{t-1}+K_{t}\nu_{t}-(1-\delta_{t-1})\imath_{t-1}-\delta_{t-1}\jmath_{t-1}$
@@
×$\qquad\times\big{(}E_{k}\boldsymbol{\nu}_{0:k-1}+K_{k}\nu_{k}+c_{k}\big{)}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
@@
$\qquad+\nu_{k}^{T}K_{k}^{T}K_{k}\nu_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
@@
×$\qquad\qquad\qquad\qquad\times\operatorname{\mathsf{p}}(\nu_{k})d\varpi_{k-1}d\nu_{k}.$
@@
$\int_{\mathbb{R}^{m}}g_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
$\leq\int_{\mathbb{R}^{m}}g^{*}_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})\big{(}\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)\big{)}^{*}d\varpi_{k-1}$
$=\int_{\mathbb{R}^{m}}g_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})\big{(}\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)\big{)}^{*}d\varpi_{k-1}$
$\leq\int_{\mathbb{R}^{m}}g_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
@@
[…]
×$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-1}=0)d\varpi_{k-1}$
[…]
×$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-1}=0)d\varpi_{k-1}$ (48)
@@
×$\qquad\qquad\qquad\qquad\times\operatorname{\mathsf{p}}(\nu_{k})d\varpi_{k-1}d\nu_{k}$
@@
×$\qquad\qquad\qquad\qquad\times\operatorname{\mathsf{p}}(\nu_{k})d\varpi_{k-1}d\nu_{k}$
$\quad=\operatorname{\mathsf{E}}_{\epsilon^{s}}\Big{[}\tilde{e}_{k}^{T}\tilde{e}_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
@@
×$\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\boldsymbol{\nu}_{0:k+1}|\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)d\boldsymbol{\nu}_{0:k+1}.$

### v2

@@ VIII Conclusions
$\qquad\qquad+\sum_{t^{\prime}=1}^{t}\operatorname{\mathsf{p}}_{\epsilon}\Big(\boldsymbol{\sigma}_{0:t^{\prime}-1}=0,\sigma_{t^{\prime}}=1\Big)=1,$ (82)
for any $t\in\mathbb{N}_{[0,N]}$. Applying the law of total expectation for the terms $\operatorname{\mathsf{E}}[\alpha_{k}u_{k}]$ and $\operatorname{\mathsf{E}}[\tilde{e}_{k}^{T}\tilde{e}_{k}]$ in $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ on a partition provided by the identity (Proof:) for t=k1t=k-1, repeating this procedure for $k\in\mathbb{N}_{[1,N]}$, and using the definition of $\Psi^{k,N}_{\epsilon}(\tilde{e}_{k})$, we can obtain (8). ∎
@@
Suppose that $(\epsilon^{o},\delta^{o})$ is a coding policy profile that belongs to the set of globally optimal solutions, which cannot be empty. We prove global optimality of the policy profile $(\epsilon^{\star},\delta^{\star})$ in the claim by showing that $\Phi(\epsilon^{\star},\delta^{\star})$ cannot be greater than $\Phi^{o}:=\Phi(\epsilon^{o},\delta^{o})$. Without loss of generality, we assume that $m_{0}=0$. Similar arguments can be made for $m_{0}\neq 0$ following a coordinate transformation. The proof consists of three steps.
@@
Step 1. We say a scheduling policy is innovation-based if, at each time $k$, it depends on $\boldsymbol{\nu}_{0:k}$ instead of $\boldsymbol{y}_{0:k}$, i.e., if it can be expressed in the form $\operatorname{\mathsf{p}}(u_{k}|\boldsymbol{\nu}_{0:k},\boldsymbol{z}^{b}_{0:k},\boldsymbol{u}_{0:k-1},\boldsymbol{\theta}_{0:k-1})$. In the first step of our proof, we show that $\Phi(\epsilon^{n},\delta^{o})=\Phi(\epsilon^{o},\delta^{o})$, where $\epsilon^{n}$ is an innovation-based scheduling policy. From the definition of the innovation, (58), and (59), we can write
$\boldsymbol{y}_{0:k}$ $=\boldsymbol{\nu}_{0:k}+G_{k}\check{\boldsymbol{x}}_{0:k-1},$
$\check{\boldsymbol{x}}_{0:k}$ $=H_{k}\boldsymbol{\nu}_{0:k},$
where $G_{k}$ and $H_{k}$ are matrices of proper dimensions. Putting these equations together, we find $\boldsymbol{y}_{0:k}=\boldsymbol{\nu}_{0:k}+G_{k}H_{k-1}\boldsymbol{\nu}_{0:k-1}$. Therefore, $\operatorname{\mathsf{p}}_{\epsilon^{o}}(u_{k}|\boldsymbol{y}_{0:k},\boldsymbol{z}^{b}_{0:k},\boldsymbol{u}_{0:k-1},\boldsymbol{\theta}_{0:k-1})$ can equivalently be written as $\operatorname{\mathsf{p}}_{\epsilon^{n}}(u_{k}|\boldsymbol{\nu}_{0:k},\boldsymbol{z}^{b}_{0:k},\boldsymbol{u}_{0:k-1},\boldsymbol{\theta}_{0:k-1})$. This establishes that $\Phi(\epsilon^{n},\delta^{o})=\Phi(\epsilon^{o},\delta^{o})$. As our subsequent analysis is valid for any values of $\boldsymbol{z}^{b}_{0:k}$ and $\boldsymbol{\theta}_{0:k-1}$, for brevity, hereafter we omit the dependency of $\epsilon^{n}$ on these variables.
@@
Step 2. Let $\mathcal{B}(r)$ be a ball of radius $r$ centred at the origin and of proper dimension. Define $\varpi_{k}:=T_{k}\boldsymbol{\nu}_{0:k}\in\mathcal{N}$ for a specific matrix $T_{k}$, and $\sigma_{k}:=u_{k}\gamma_{k}$, $l_{\epsilon^{s}}(\varpi_{k}):=(\lambda+\lambda^{c}\operatorname{\mathsf{p}}_{\epsilon^{s}}(u_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0))\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$, and $l_{\epsilon^{n}}(\varpi_{k}):=(\lambda+\lambda^{c}\operatorname{\mathsf{p}}_{\epsilon^{n}}(u_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0))\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$. In the second step of our proof, we show that $\Phi(\epsilon^{s},\delta^{o})\leq\Phi(\epsilon^{n},\delta^{o})$, where $\epsilon^{s}$ is a special form of $\epsilon^{n}$ that is symmetric with respect to $\boldsymbol{\nu}_{0:k}$ at each time $k$. Let us construct $\epsilon^{s}$ such that the following condition is satisfied:
$\int_{\mathcal{B}(r)}l_{\epsilon^{s}}(\varpi_{k})d\varpi_{k}=\int_{\mathcal{B}(r)}l^{*}_{\epsilon^{n}}(\varpi_{k})d\varpi_{k},$ (83)
@@
for all $r\geq 0$ with $l_{\epsilon^{s}}(\varpi_{k})$ as a radially symmetric function of $\varpi_{k}$. We know that $\sigma_{k}=0$ either when $u_{k}=1\;\wedge\;\gamma_{k}=0$ or when $u_{k}=0$ regardless of $\gamma_{k}$. Accordingly, we can write
@@
$\ =\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big(u_{k}=1\big|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0)$
$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big(u_{k}\gamma_{k}=0\big|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0,u_{k}=1\Big)$
$\qquad+\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big(u_{k}=0\big|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big)$
$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big(u_{k}\gamma_{k}=0\big|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0,u_{k}=0\Big)$
$\ =\lambda\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big(u_{k}=1\big|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big)$
$\qquad+\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big(u_{k}=0\big|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big)$
$\ =\lambda+\lambda^{c}\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big(u_{k}=0\big|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big),$
where in the first equality we used the law of total probability and in the second equality the fact that $\gamma_{k}$ is independent of $\varpi_{k}$, $\boldsymbol{\sigma}_{0:k-1}$, and $u_{k}$. Define $l^{\prime}_{\epsilon^{s}}(\varpi_{k})=\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0)\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$ and $l^{\prime}_{\epsilon^{s}}(\varpi_{k})=\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0)\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$. Therefore, the above condition can be written in the following equivalent form:
$\int_{\mathcal{B}(r)}l^{\prime}_{\epsilon^{s}}(\varpi_{k})d\varpi_{k}=\int_{\mathcal{B}(r)}{l^{\prime}}^{*}_{\epsilon^{n}}(\varpi_{k})d\varpi_{k},$ (84)
for all $r\geq 0$ with $l^{\prime}_{\epsilon^{s}}(\varpi_{k})$ as a radially symmetric function of $\varpi_{k}$. By Lemma 8, to prove that $\Phi(\epsilon^{s},\delta^{o})\leq\Phi(\epsilon^{n},\delta^{o})$, it suffices to prove that $\Psi^{M}_{\epsilon^{s}}(\tilde{e}_{0})\leq\Psi^{M}_{\epsilon^{n}}(\tilde{e}_{0})$ for any $M\in\mathbb{N}_{[0,N]}$ and for any $\tilde{e}_{0}$. Observe that $\tilde{e}_{0}=K_{0}\nu_{0}$ is the same under both $\epsilon^{n}$ and $\epsilon^{s}$, and that $u_{0}$ has no effect on remote estimation when the time horizon is 0. Hence, the claim holds for the time horizon 0. We assume that the claim also holds for all time horizons from $1$ to M1M-1, and will show that the terms in $\Psi^{M}_{\epsilon^{n}}(\tilde{e}_{0})$ are not less than those in $\Psi^{M}_{\epsilon^{s}}(\tilde{e}_{0})$.
@@
$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
@@
$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
$\ =\operatorname{\mathsf{p}}_{\epsilon^{s}}\Big(\sigma_{k-1}=0\big|\boldsymbol{\sigma}_{0:k-2}=0\Big),$
where the second equality comes from (84) when $r\to\infty$, following the fact that $\int_{\mathbb{R}^{m}}{l^{\prime}}^{*}_{\epsilon^{n}}(\varpi_{k})d\varpi_{k}=\int_{\mathbb{R}^{m}}l^{\prime}_{\epsilon^{n}}(\varpi_{k})d\varpi_{k}$. This also implies that $\operatorname{\mathsf{p}}_{\epsilon^{n}}(\boldsymbol{\sigma}_{0:k-1}=0)=\operatorname{\mathsf{p}}_{\epsilon^{s}}(\boldsymbol{\sigma}_{0:k-1}=0)$ and that $\operatorname{\mathsf{p}}_{\epsilon^{n}}(\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)=\operatorname{\mathsf{p}}_{\epsilon^{s}}(\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)$.
@@
$\operatorname{\mathsf{E}}_{\epsilon^{n}}\Big[$ $u_{k}\big|\boldsymbol{\sigma}_{0:k-1}=0\Big]$
$=1-\operatorname{\mathsf{p}}_{\epsilon^{n}}(u_{k}=0|\boldsymbol{\sigma}_{0:k-1}=0)$
@@
$=1-\operatorname{\mathsf{p}}_{\epsilon^{s}}(u_{k}=0|\boldsymbol{\sigma}_{0:k-1}=0)$
$=\operatorname{\mathsf{E}}_{\epsilon^{s}}\Big[u_{k}\big|\boldsymbol{\sigma}_{0:k-1}=0\Big].$
@@
$\tilde{e}_{t}$ $=A_{t-1}\tilde{e}_{t-1}+K_{t}\nu_{t}-(1-u_{t-1})\imath_{t-1}-u_{t-1}\jmath_{t-1},$
@@
$\qquad\times\big(E_{k}\boldsymbol{\nu}_{0:k-1}+K_{k}\nu_{k}+c_{k}\big)\big|\boldsymbol{\sigma}_{0:k-1}=0\Big]$
@@
$\qquad+\nu_{k}^{T}K_{k}^{T}K_{k}\nu_{k}\big|\boldsymbol{\sigma}_{0:k-1}=0\Big],$
@@
$\qquad\qquad\qquad\qquad\times\operatorname{\mathsf{p}}(\nu_{k})d\varpi_{k-1}d\nu_{k}.$
@@
$\int_{\mathbb{R}^{m}}g_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
$\leq\int_{\mathbb{R}^{m}}g^{*}_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})\big(\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)\big)^{*}d\varpi_{k-1}$
$=\int_{\mathbb{R}^{m}}g_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})\big(\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)\big)^{*}d\varpi_{k-1}$
$\leq\int_{\mathbb{R}^{m}}g_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1},$
@@
$\int_{\mathbb{R}^{m}}\underset{z}{\min}\Big\{z,f_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})\Big\}$
$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-1}=0)d\varpi_{k-1}$
$\geq\int_{\mathbb{R}^{m}}\underset{z}{\min}\Big\{z,f_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})\Big\}$
$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-1}=0)d\varpi_{k-1},$ (85)
@@
$\qquad\qquad\qquad\qquad\times\operatorname{\mathsf{p}}(\nu_{k})d\varpi_{k-1}d\nu_{k}$
@@
$\qquad\qquad\qquad\qquad\times\operatorname{\mathsf{p}}(\nu_{k})d\varpi_{k-1}d\nu_{k}$
$\quad=\operatorname{\mathsf{E}}_{\epsilon^{s}}\Big[\tilde{e}_{k}^{T}\tilde{e}_{k}\big|\boldsymbol{\sigma}_{0:k-1}=0\Big],$
@@
$\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\boldsymbol{\nu}_{0:k+1}|\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)d\boldsymbol{\nu}_{0:k+1}.$